OpenStudy (anonymous):

What is the 9th term of the geometric sequence where a1 = –6 and a6 = –6,144?

OpenStudy (anonymous):

a6 = a1 * q^5 -6.144 = -6 * q^5 q^5 = + 1024 q = sqrt 5 of 1024 q = sqrt 5 of 2^(10) q = 2^ ( 10/5) =2^2 = 4 <--- important a9 = a1 * q^8 a9 = -6 * 4^8

OpenStudy (anonymous):

-393216=T9 a=-6 t6=-6144 -6144=a.r^n-1 -6144=-6 .r^5 1024=r^5 r=common ratio=4 t9=a.r^8 -393216

OpenStudy (anonymous):

in gp for a next term tn=a*r^n-1 in Ap=a+(n-1)d